### v1

@@ 1 Introduction
QCD jets are collimated sprays of hadrons with fluctuating particle numbers, and their multiplicity distributions, that is, the probability distribution $P(n)$ for producing $n$ particles, constitute an essential feature. Studying these distributions provides valuable insights into the dynamics of Quantum Chromodynamics (QCD), particularly the mechanisms of parton branching and hadronization in the high-virtuality regime. An expected simplification of multiplicity distributions in high-energy QCD jets arises from the concept of Koba–Nielsen–Olesen (KNO) scaling Koba:1972ng . This scaling hypothesis states that at asymptotically high energy, the probability distribution of final-state multiplicities $P(n)$ becomes self-similar when rescaled by the mean multiplicity $\bar{n}$, such that $\bar{n}P(n)$ depends only on the scaling variable $x=n/\bar{n}$ and approaches a universal function $\Psi(x)$, that is,
@@
with $\int dx~\Psi(x)=1$. It was originally proposed on general grounds of scale invariance in quantum field theory Polyakov:1970lyy ; Feynman:1969ej , and found a natural interpretation within the framework of QCD as a consequence of the cascade structure of parton showers in QCD jets Konishi:1979cb ; Bassetto:1979nt ; Dokshitzer:1982ia ; Malaza:1984vv ; Bassetto:1987fq ; Dokshitzer:1991wu ; Dokshitzer:1993dc ; Vertesi:2020utz ; Germano:2024ier ; Duan:2025ngi ; Dokshitzer:2025owq ; Dokshitzer:2025fky .
@@
Multiplicity distributions are expected to exhibit universality across different high-energy collision systems, including $e^{+}e^{-}$, Deep Inelastic Scattering (DIS), and $pp$ collisions, when dominated by high-energy jets. In particular, measurements of hadronic final states in $e^{+}e^{-}$ annihilation over a wide energy range, from PETRA to LEP TASSO:1983cre ; TASSO:1989orr ; DELPHI:1990ohs ; DELPHI:1991qnt , showed that charged-particle multiplicity distributions approximately collapse onto a universal KNO curve. Nonetheless, deviations from exact KNO scaling are expected, arising from the superposition of jets with different flavors and topologies ALEPH:1995qic ; Giovannini:1997ce , as well as from the running of the strong coupling Dokshitzer:2025owq ; Dokshitzer:2025fky . Similar approximate KNO behavior is observed in DIS at HERA H1:2020zpd , where both theoretical predictions Liu:2022bru ; Liu:2023eve and experimental data support scaling in the current hemisphere at large momentum transfer. Early low-energy $pp$ collisions, from bubble chamber experiments to ISR data, also reported approximate scaling of multiplicity distributions Slattery:1972ni ; Ames-Bologna-CERN-Dortmund-Heidelberg-Warsaw:1983cqw . In high-energy hadronic collisions, however, LHC data UA5:1985kkp ; UA5:1988gup ; CMS:2010qvf ; Grosse-Oetringhaus:2009eis ; ALICE:2017pcy show that multiplicity distributions broaden with increasing energy and no longer follow KNO scaling. This breakdown is not surprising, since contributions from mixed event topologies of hard processes, such as quark- and gluon-initiated jets Duan:2025ngi ; Dokshitzer:2025owq , as well as soft processes, can lead to scaling violation (for recent discussions, see ref. Martins-Fontes:2025iee and references therein). KNO scaling is suggested to hold within QCD jets in $pp$ collisions Vertesi:2020utz ; Germano:2024ier , while such universality is anticipated to reemerge when focusing specifically on high-energy QCD jets with proper quark–gluon discrimination Duan:2025ngi ; Dokshitzer:2025owq .
@@
The KNO scaling functions are different for quark- and gluon-initiated jets, referred to as quark and gluon jets below, and have been calculated in perturbative QCD Dokshitzer:1982ia ; Bassetto:1987fq ; Dokshitzer:1991wu ; Dokshitzer:1993dc ; Duan:2025ngi ; Dokshitzer:2025owq ; Dokshitzer:2025fky . In the double logarithmic approximation (DLA), where the dominant soft and collinear emissions are resummed with strong ordering in both energy and angle, the forms of the scaling functions $\Psi(x)$ for gluon and quark jets can be evaluated explicitly Dokshitzer:1982ia ; Bassetto:1987fq ; Dokshitzer:1991wu . These scaling functions are exact in the asymptotic high-energy (high-virtuality) limit, and KNO scaling within DLA has recently been confirmed to hold to a good approximation by numerically solving the multiplicity distributions in QCD jets for a broad jet $p_{T}$ range relevant to LHC energies Duan:2025ngi .
@@
However, the DLA KNO scaling functions do not provide a reliable description of the measured multiplicity distributions Malaza:1984vv ; Dokshitzer:1993dc ; Duan:2025ngi ; Dokshitzer:2025owq ; Dokshitzer:2025fky . To remedy this, the DLA for gluon jets was first improved by enforcing energy conservation in the parton cascade Dokshitzer:1993dc . This modified double logarithmic approximation (MDLA) has recently been generalized to jets initiated by any source, including quarks and antiquarks, achieving notable phenomenological success Dokshitzer:2025owq ; Dokshitzer:2025fky . Meanwhile, the KNO scaling functions have also been extracted using PYTHIA Vertesi:2020utz ; Duan:2025ngi , showing good agreement with LHC measurements within experimental uncertainties ATLAS:2019rqw ; Duan:2025ngi . One of the main goals of this work is to compare quark and gluon KNO scaling functions obtained within MDLA and PYTHIA, and to explore their phenomenological implications.
@@
Recent advances in jet physics and jet substructure at the LHC Marzani:2019hun ; Larkoski:2024uoc provide an unprecedented opportunity to investigate multiplicity fluctuations in QCD jets at very high transverse momenta, both theoretically and experimentally. A variety of methods for quark–gluon discrimination have been proposed Gallicchio:2011xq ; Gallicchio:2012ez ; Larkoski:2013eya ; Larkoski:2014pca ; Gras:2017jty ; Metodiev:2017vrx ; Metodiev:2018ftz ; Komiske:2018vkc ; Dreyer:2021hhr and applied in LHC analyses CMS:2013kfa ; ATLAS:2014vax ; ATLAS:2016wzt ; ATLAS:2019rqw . If proven effective for multiplicity distributions, these techniques would enable direct experimental tests of the universality of the KNO scaling functions for quark and gluon jets, as illustrated using energy–correlation functions Larkoski:2013eya in ref. Duan:2025ngi . Another central motivation of this paper is to investigate KNO scaling in QCD jets using jet topics Metodiev:2018ftz ; Komiske:2018vkc , which is well suited to extracting multiplicity distributions and can be implemented without additional theoretical input.
@@
The rest of this paper is structured as follows. In section 2, we start with a review of the generating function method for multiplicity distributions within QCD jets, as well as a summary of the DLA results for parton multiplicity distributions and mean multiplicities. We then calculate the KNO scaling function for both quark and gluon jets within the MDLA, i.e., by incorporating energy conservation into the DLA, following refs. Dokshitzer:1982ia ; Bassetto:1987fq ; Dokshitzer:1993dc , as an alternative to the recent calculations in refs. Dokshitzer:2025owq ; Dokshitzer:2025fky . In section 3, we present a phenomenological study of inclusive charged-particle multiplicity distributions of two leading jets in $pp$ collisions at $\sqrt{s}=13$ TeV, combining the MDLA KNO scaling functions with leading-order cross sections and the calculations of mean multiplicities, and compare our theoretical results with ATLAS data from ref. ATLAS:2019rqw . As a consistency check, we also directly compare the MDLA KNO scaling functions for quark and gluon jets with those extracted from ATLAS measurements using topic modeling Metodiev:2018ftz ; Komiske:2018vkc , which is presented in section 4. We conclude in section 5 with a summary of results and outlook for future directions. Finally, for completeness, we collect the next-to-next-to-next-to-leading-order (N3LO) results for mean parton multiplicities in QCD jets from refs. Dremin:1999ji ; Capella:1999ms in Appendix A, which are used in our phenomenological studies in section 3.
@@ 2 Multiplicity distributions and KNO scaling in QCD jets
In this section, we begin with a brief review—for the sole purpose of completeness in our subsequent discussion—of the method of generating functions (GFs) Konishi:1979cb ; Bassetto:1979nt ; Dokshitzer:1982ia ; Dokshitzer:1991wu ; Ellis:1996mzs ; Dremin:2000ep and the DLA results for parton multiplicity distributions within QCD jets Dokshitzer:1982xr ; Dokshitzer:1982ia ; Dokshitzer:1982fh ; Duan:2025ngi . We then evaluate the KNO scaling functions within MDLA for both quark and gluon jets by building on the treatment of gluon jets presented in ref. Dokshitzer:1993dc , providing an alternative to the recent calculations in refs. Dokshitzer:2025owq ; Dokshitzer:2025fky .
@@ 2.1 Review of the generating function method
The GF method serves as a powerful and convenient tool for describing the evolution of QCD jets produced in high-energy hard scatterings. To support the forthcoming discussion, we outline the basics of the method and summarize the DLA results of parton multiplicities within QCD jets here. For a more comprehensive discussion, interested readers are referred to refs. Dokshitzer:1991wu ; Ellis:1996mzs ; Dremin:2000ep .
@@ 2.1.1 The definition of generating functions
Analogous to the integral form of the DGLAP equation, the GFs at next-to-leading logarithmic (NLL) order obey Ellis:1996mzs
@@
Equivalently, by differentiating eq. (2.1.1) with respect to $Q$, one obtains, at $O(\alpha_{s})$ Dremin:2000ep ,
@@
The first term on the right-hand side of eq. (2.1.1), involving the form factor $\Delta_{a}$, corresponds to the process in which the initial parton $a$ remains unchanged, while the second term accounts for the splitting process $a\to bc$. When coherent branching (i.e., angular ordering) Dokshitzer:1987nm ; Dokshitzer:1991wu is incorporated into the second term, the formalism properly accounts for both soft and collinear enhancements. Denote the momenta of the daughter parton $b$ and its parent $a$ by $k$ and $p$, respectively. In this case, the evolution variable is most conveniently defined as Ellis:1996mzs
@@ 2.1.2 Parton multiplicities in QCD jets within DLA
Focusing solely on soft and collinear logarithmic contributions, the splitting function simplifies to $\hat{P}_{a\to ga}(z)\approx 2C_{a}/z$. And the condition for coherent branching can be expressed as Dokshitzer:1991wu
@@
For hadron multiplicities, the GF necessarily includes non-perturbative contributions. In this section, we instead focus on parton multiplicities at the reference virtuality $Q_{0}$, where the system contains a single parton. This sets the initial condition: $Z_{a}(u,Q_{0})=u$.111 Note that in refs. Dokshitzer:1982xr ; Dokshitzer:1982ia ; Dokshitzer:1982fh , the GFs account only for the radiated soft gluons with $Z_{a}(u,Q_{0})=1$. In contrast, our convention follows ref. Dokshitzer:1991wu , where the original leading parton is counted in $P_{n}$. Restricting to the DLA phase space in eq. (9), in terms of $k_{\perp}=z\bar{Q}$ and $z$, one obtains Dokshitzer:1991wu
[…]
where $y\equiv\ln(Q/Q_{0})$ and $\bar{y}\equiv\ln(k_{\perp}/Q_{0})$. From the GFs given in eq. (12), we obtain the following recursive relation between $P_{a}(n)$’s Duan:2025ngi :
[…]
@@
$\frac{\partial^{2}}{\partial y^{2}}\bar{n}_{g}(Q)$ $=\gamma_{0}^{2}\bar{n}_{g}$ (16)
with the boundary conditions $\bar{n}_{g}^{\prime}(0)=0$ and $\bar{n}_{g}(0)=1$. This second-order ordinary differential equation admits the analytic solution: Dokshitzer:1982fh ; Dokshitzer:1982xr ; Dokshitzer:1991wu
@@
where $z_{1}\equiv A\sqrt{Y}$ and $z_{2}\equiv A\sqrt{\lambda}$, with $Y\equiv y+\lambda=\ln({Q}/{\Lambda})$, $A\equiv\sqrt{{16N_{c}}/{\beta_{0}}}$ and $\beta_{0}=({11}N_{c}-{2}n_{f})/{3}$. Alternatively, given $P_{a}(n,Q)$ numerically evaluated according to eq. (2.1.2), the mean parton multiplicities in quark and gluon jets can be obtained via:
@@
The agreement of these two approaches has been verified in ref. Duan:2025ngi .

### v2

@@ 1 Introduction
QCD jets are collimated sprays of hadrons with fluctuating particle numbers, and their multiplicity distributions, that is, the probability distribution $P(n)$ for producing $n$ particles, constitute an essential feature. Studying these distributions provides valuable insights into the dynamics of Quantum Chromodynamics (QCD), particularly the mechanisms of parton branching and hadronization in the high-virtuality regime. An expected simplification of multiplicity distributions in high-energy QCD jets arises from the concept of Koba–Nielsen–Olesen (KNO) scaling Koba et al. (1972). This scaling hypothesis states that at asymptotically high energy, the probability distribution of final-state multiplicities $P(n)$ becomes self-similar when rescaled by the mean multiplicity $\bar{n}$, such that $\bar{n}P(n)$ depends only on the scaling variable $x=n/\bar{n}$ and approaches a universal function $\Psi(x)$, that is,
@@
with $\int dx~\Psi(x)=1$. It was originally proposed on general grounds of scale invariance in quantum field theory Polyakov (1970); Feynman (1969), and found a natural interpretation within the framework of QCD as a consequence of the cascade structure of parton showers in QCD jets Konishi et al. (1979); Bassetto et al. (1980); Dokshitzer et al. (1983); Malaza and Webber (1984); Bassetto (1988); Dokshitzer et al. (1991); Dokshitzer (1993); Vertesi et al. (2021); Germano et al. (2024); Duan et al. (2025); Dokshitzer and Webber (2025a, b).
@@
Multiplicity distributions are expected to exhibit universality across different high-energy collision systems, including $e^{+}e^{-}$, Deep Inelastic Scattering (DIS), and $pp$ collisions, when dominated by high-energy jets. In particular, measurements of hadronic final states in $e^{+}e^{-}$ annihilation over a wide energy range, from PETRA to LEP Althoff and others (1984); Braunschweig and others (1989); Abreu and others (1991b, a), showed that charged-particle multiplicity distributions approximately collapse onto a universal KNO curve. Nonetheless, deviations from exact KNO scaling are expected, arising from the superposition of jets with different flavors and topologies Buskulic and others (1995); Giovannini and Ugoccioni (1999), as well as from the running of the strong coupling Dokshitzer and Webber (2025a, b). Similar approximate KNO behavior is observed in DIS at HERA Andreev and others (2021), where both theoretical predictions Liu et al. (2023a, b) and experimental data support scaling in the current hemisphere at large momentum transfer. Early low-energy $pp$ collisions, from bubble chamber experiments to ISR data, also reported approximate scaling of multiplicity distributions Slattery (1972); Breakstone and others (1984). In high-energy hadronic collisions, however, LHC data Alner and others (1985); Ansorge and others (1989); Khachatryan and others (2011); Grosse-Oetringhaus and Reygers (2010); Acharya and others (2017) show that multiplicity distributions broaden with increasing energy and no longer follow KNO scaling. This breakdown is not surprising, since contributions from mixed event topologies of hard processes, such as quark- and gluon-initiated jets Duan et al. (2025); Dokshitzer and Webber (2025a), as well as soft processes, can lead to scaling violation (for recent discussions, see ref. Martins-Fontes and Navarra (2025) and references therein). KNO scaling is suggested to hold within QCD jets in $pp$ collisions Vertesi et al. (2021); Germano et al. (2024), while such universality is anticipated to reemerge when focusing specifically on high-energy QCD jets with proper quark–gluon discrimination Duan et al. (2025); Dokshitzer and Webber (2025a).
@@
The KNO scaling functions are different for quark- and gluon-initiated jets, referred to as quark and gluon jets below, and have been calculated in perturbative QCD Dokshitzer et al. (1983); Bassetto (1988); Dokshitzer et al. (1991); Dokshitzer (1993); Duan et al. (2025); Dokshitzer and Webber (2025a, b). In the double logarithmic approximation (DLA), where the dominant soft and collinear emissions are resummed with strong ordering in both energy and angle, the forms of the scaling functions $\Psi(x)$ for gluon and quark jets can be evaluated explicitly Dokshitzer et al. (1983); Bassetto (1988); Dokshitzer et al. (1991). These scaling functions are exact in the asymptotic high-energy (high-virtuality) limit, and KNO scaling within DLA has recently been confirmed to hold to a good approximation by numerically solving the multiplicity distributions in QCD jets for a broad jet $p_{T}$ range relevant to LHC energies Duan et al. (2025).
@@
However, the DLA KNO scaling functions do not provide a reliable description of the measured multiplicity distributions Malaza and Webber (1984); Dokshitzer (1993); Duan et al. (2025); Dokshitzer and Webber (2025a, b). To remedy this, the DLA for gluon jets was first improved by enforcing energy conservation in the parton cascade Dokshitzer (1993). This modified double logarithmic approximation (MDLA) has recently been generalized to jets initiated by any source, including quarks and antiquarks, achieving notable phenomenological success Dokshitzer and Webber (2025a, b). Meanwhile, the KNO scaling functions have also been extracted using PYTHIA Vertesi et al. (2021); Duan et al. (2025), showing good agreement with LHC measurements within experimental uncertainties Aad and others (2019); Duan et al. (2025). One of the main goals of this work is to compare quark and gluon KNO scaling functions obtained within MDLA and PYTHIA, and to explore their phenomenological implications.
@@
Recent advances in jet physics and jet substructure at the LHC Marzani et al. (2019); Larkoski (2024) provide an unprecedented opportunity to investigate multiplicity fluctuations in QCD jets at very high transverse momenta, both theoretically and experimentally. A variety of methods for quark–gluon discrimination have been proposed Gallicchio and Schwartz (2011, 2013); Larkoski et al. (2013, 2014); Gras et al. (2017); Metodiev et al. (2017); Metodiev and Thaler (2018); Komiske et al. (2018); Dreyer et al. (2022) and applied in LHC analyses 62; G. Aad et al. (2014); 23; G. Aad et al. (2019). If proven effective for multiplicity distributions, these techniques would enable direct experimental tests of the universality of the KNO scaling functions for quark and gluon jets, as illustrated using energy–correlation functions Larkoski et al. (2013) in ref. Duan et al. (2025). Another central motivation of this paper is to investigate KNO scaling in QCD jets using jet topics Metodiev and Thaler (2018); Komiske et al. (2018), which is well suited to extracting multiplicity distributions and can be implemented without additional theoretical input.
@@
The rest of this paper is structured as follows. In section 2, we start with a review of the generating function method for multiplicity distributions within QCD jets, as well as a summary of the DLA results for parton multiplicity distributions and mean multiplicities. We then calculate the KNO scaling function for both quark and gluon jets within the MDLA, i.e., by incorporating energy conservation into the DLA, following refs. Dokshitzer et al. (1983); Bassetto (1988); Dokshitzer (1993), as an alternative to the recent calculations in refs. Dokshitzer and Webber (2025a, b). In section 3, we present a phenomenological study of inclusive charged-particle multiplicity distributions of two leading jets in $pp$ collisions at $\sqrt{s}=13$ TeV, combining the MDLA KNO scaling functions with leading-order cross sections and the calculations of mean multiplicities, and compare our theoretical results with ATLAS data from ref. Aad and others (2019). As a consistency check, we also directly compare the MDLA KNO scaling functions for quark and gluon jets with those extracted from ATLAS measurements using topic modeling Metodiev and Thaler (2018); Komiske et al. (2018), which is presented in section 4. We conclude in section 5 with a summary of results and outlook for future directions. Finally, for completeness, we collect the next-to-next-to-next-to-leading-order (N3LO) results for mean parton multiplicities in QCD jets from refs. Dremin and Gary (1999); Capella et al. (2000) in Appendix A, which are used in our phenomenological studies in section 3.
@@ 2 Multiplicity distributions and KNO scaling in QCD jets
In this section, we begin with a brief review, for the sole purpose of completeness in our subsequent discussion, of the method of generating functions (GFs) Konishi et al. (1979); Bassetto et al. (1980); Dokshitzer et al. (1983, 1991); Ellis et al. (2011); Dremin and Gary (2001) and the DLA results for parton multiplicity distributions within QCD jets Dokshitzer et al. (1982b, 1983, a); Duan et al. (2025). We then evaluate the KNO scaling functions within MDLA for both quark and gluon jets by building on the treatment of gluon jets presented in ref. Dokshitzer (1993), providing an alternative to the recent calculations in refs. Dokshitzer and Webber (2025a, b).
@@ 2.1 Review of the generating function method
The GF method serves as a powerful and convenient tool for describing the evolution of QCD jets produced in high-energy hard scatterings. To support the forthcoming discussion, we outline the basics of the method and summarize the DLA results of parton multiplicities within QCD jets here. For a more comprehensive discussion, interested readers are referred to refs. Dokshitzer et al. (1991); Ellis et al. (2011); Dremin and Gary (2001).
@@ 2.1.1 The definition of generating functions
Analogous to the integral form of the DGLAP equation, the GFs at next-to-leading logarithmic (NLL) order obey Ellis et al. (2011)
@@
Equivalently, by differentiating eq. (2.1.1) with respect to $Q$, one obtains, at $O(\alpha_{s})$ Dremin and Gary (2001),
@@
The first term on the right-hand side of eq. (2.1.1), involving the form factor $\Delta_{a}$, corresponds to the process in which the initial parton $a$ remains unchanged, while the second term accounts for the splitting process $a\to bc$. When coherent branching (i.e., angular ordering) Dokshitzer et al. (1988, 1991) is incorporated into the second term, the formalism properly accounts for both soft and collinear enhancements. Denote the momenta of the daughter parton $b$ and its parent $a$ by $k$ and $p$, respectively. In this case, the evolution variable is most conveniently defined as Ellis et al. (2011)
@@ 2.1.2 Parton multiplicities in QCD jets within DLA
Focusing solely on soft and collinear logarithmic contributions, the splitting function simplifies to $\hat{P}_{a\to ga}(z)\approx 2C_{a}/z$. And the condition for coherent branching can be expressed as Dokshitzer et al. (1991)
@@
For hadron multiplicities, the GF necessarily includes non-perturbative contributions. In this section, we instead focus on parton multiplicities at the reference virtuality $Q_{0}$, where the system contains a single parton. This sets the initial condition: $Z_{a}(u,Q_{0})=u$. [^1] Note that in refs. Dokshitzer et al. (1982b, 1983, a), the GFs account only for the radiated soft gluons with $Z_{a}(u,Q_{0})=1$. In contrast, our convention follows ref. Dokshitzer et al. (1991), where the original leading parton is counted in $P_{a}(n)$. Restricting to the DLA phase space in eq. (9), in terms of $k_{\perp}=z\bar{Q}$ and $z$, one obtains Dokshitzer et al. (1991)
$Z_{a}(u,y)=u\exp\left\{c_{a}\int_{0}^{y}d\bar{y}~(y-\bar{y})\gamma_{0}^{2}[Z_{g}(u,\bar{y})-1]\right\},$ (12)
where $y\equiv\ln(Q/Q_{0})$ and $\bar{y}\equiv\ln(k_{\perp}/Q_{0})$. From the GFs given in eq. (12), we obtain the following recursive relation between $P_{a}(n)$’s Duan et al. (2025):
$P_{a}(1,Q)$ $=\exp\left\{-c_{a}\int_{0}^{y}d\bar{y}~(y-\bar{y})\gamma_{0}^{2}\right\},$
$P_{a}(n+1,Q)$ $=c_{a}\sum_{k=1}^{n}\frac{k}{n}P_{a}(n+1-k,Q)\int_{0}^{y}d\bar{y}~(y-\bar{y})\gamma_{0}^{2}P_{g}(k,\bar{y}).$ (13)
@@
$\frac{\partial^{2}}{\partial y^{2}}\bar{n}_{g}$ $=\gamma_{0}^{2}\bar{n}_{g}$ (16)
with the boundary conditions $\bar{n}_{g}^{\prime}(0)=0$ and $\bar{n}_{g}(0)=1$. This second-order ordinary differential equation admits the analytic solution: Dokshitzer et al. (1982a, b, 1991)
@@
where $z_{1}\equiv A\sqrt{y+\lambda}$ and $z_{2}\equiv A\sqrt{\lambda}$, with $\lambda\equiv\ln(Q_{0}/\Lambda)$, $A\equiv\sqrt{{16N_{c}}/{\beta_{0}}}$, and $\beta_{0}=({11}N_{c}-{2}n_{f})/{3}$. Alternatively, given $P_{a}(n,Q)$ numerically evaluated according to eq. (2.1.2), the mean parton multiplicities in quark and gluon jets can be obtained via:
@@
The agreement of these two approaches has been verified in ref. Duan et al. (2025).